# Inference-Time Decomposition of Activations (ITDA): A Scalable Approach to Interpreting Large Language Models

### v1

@@ Abstract
Sparse autoencoders (SAEs) are a popular method for decomposing Large Langage Models (LLM) activations into interpretable latents. However, due to their substantial training cost, most academic research uses open-source SAEs which are only available for a restricted set of models of up to 27B parameters. SAE latents are also learned from a dataset of activations, which means they do not transfer between models. Motivated by relative representation similarity measures, we introduce Inference-Time Decomposition of Activations (ITDA) models, an alternative method for decomposing language model activations. To train an ITDA, we greedily construct a dictionary of language model activations on a dataset of prompts, selecting those activations which were worst approximated by matching pursuit on the existing dictionary. ITDAs can be trained in just 1% of the time required for SAEs, using 1% of the data. This allowed us to train ITDAs on Llama-3.1 70B and 405B on a single consumer GPU. ITDAs can achieve similar reconstruction performance to SAEs on some target LLMs, but generally incur a performance penalty. However, ITDA dictionaries enable cross-model comparisons, and a simple Jaccard similarity index on ITDA dictionaries outperforms existing methods like CKA, SVCCA, and relative representation similarity metrics. ITDAs provide a cheap alternative to SAEs where computational resources are limited, or when cross model comparisons are necessary. Code available at github.com/pleask/itda.
@@ 1 Introduction
Furthermore, comparing SAEs trained on different models is challenging: their parameters are learned from model activations, which vary for the same input between models even of the same family. Lan et al. (2024) proposes that SAEs reveal universal feature spaces across LLMs by using centered kernel alignment (Kornblith et al., 2019) to compare the decoders of SAEs. The problem of aligning the representation spaces of different LLMs is also fundamental to absolute measures of representation similarity such as SVCCA (Raghu et al., 2017) and CKA (Kornblith et al., 2019), which is addressed with relative representation measures by Moschella et al. (2022), who use the cosine similarity of representations with respect to the representations of a fixed random set of anchor inputs in different models.
@@ 2.2 SAEs for Mechanistic Interpretability
SAEs have been demonstrated to recover sparse, monosemantic, and interpretable features from language model representations (Bricken et al., 2023; Cunningham et al., 2023; Templeton, 2024; Gao et al., 2025; Rajamanoharan et al., 2025, 2024), however their application to mechanistic interpretability is nascent. After training, researchers often interpret the meaning of SAE latents by examining the dataset examples on which they are active, either through manual inspection using features dashboards (Bricken et al., 2023) or automated interpretability techniques (Gao et al., 2025). SAEs have been used for circuit analysis (Marks et al., 2025) in the vein of (Olah et al., 2020; Olsson et al., 2022); to study the role of attention heads in GPT-2 (Kissane et al., 2024a); and to replicate the identification of a circuit for indirect object identification in GPT-2 (Makelov et al., 2025). Transcoders, a variant of SAEs, have been used to simplify circuit analysis and applied to the greater-than circuit in GPT-2 (Dunefsky et al., 2024). Whilst the application of SAEs to mechanistic interpretability is supported by qualitative and quantitative evidence, their usefulness is highly dependent on hyperparameterisation (Leask et al., 2025). (Nanda et al., 2024) propose the use of inference-time optimization on sparse autoencoder decoder matrices to quickly evaluate and compare them.

### v2

@@ Abstract
Sparse Autoencoders (SAEs) are a popular method for decomposing Large Language Model (LLM) activations into interpretable latents, however they have a substantial training cost and SAEs learned on different models are not directly comparable. Motivated by relative representation similarity measures, we introduce Inference-Time Decomposition of Activation models (ITDAs). ITDAs are constructed by greedily sampling activations into a dictionary based on an error threshold on their matching pursuit reconstruction. ITDAs can be trained in 1% of the time of SAEs, allowing us to cheaply train them on Llama-3.1 70B and 405B. ITDA dictionaries also enable cross-model comparisons, and outperform existing methods like CKA, SVCCA, and a relative representation method on a benchmark of representation similarity. Code available at github.com/pleask/itda.
@@ 1 Introduction
Furthermore, comparing SAEs trained on different models is challenging: their parameters are learned from model activations, which vary for the same input between models even of the same family. Lan et al. (2024) proposes that SAEs reveal universal feature spaces across LLMs by using SVCCA (Raghu et al., 2017) to compare the decoders of SAEs. The problem of aligning the representation spaces of different LLMs is also fundamental to absolute measures of representation similarity such as SVCCA (Raghu et al., 2017) and CKA (Kornblith et al., 2019), which is addressed with relative representation measures by Moschella et al. (2022), who use the cosine similarity of representations with respect to the representations of a fixed random set of anchor inputs in different models.
@@ 2.2 SAEs for Mechanistic Interpretability
SAEs have been demonstrated to recover sparse, monosemantic, and interpretable features from language model representations (Bricken et al., 2023; Cunningham et al., 2023; Templeton, 2024; Gao et al., 2025; Rajamanoharan et al., 2025, 2024), however their application to mechanistic interpretability is nascent. After training, researchers often interpret the meaning of SAE latents by examining the dataset examples on which they are active, either through manual inspection using features dashboards (Bricken et al., 2023) or automated interpretability techniques (Gao et al., 2025; Paulo et al., 2024). SAEs have been used for circuit analysis (Marks et al., 2025) in the vein of (Olah et al., 2020; Olsson et al., 2022); to study the role of attention heads in GPT-2 (Kissane et al., 2024a); and to replicate the identification of a circuit for indirect object identification in GPT-2 (Makelov et al., 2025). Transcoders, a variant of SAEs, have been used to simplify circuit analysis and applied to the greater-than circuit in GPT-2 (Dunefsky et al., 2024). Whilst the application of SAEs to mechanistic interpretability is supported by qualitative and quantitative evidence, their usefulness is highly dependent on hyperparameterisation (Leask et al., 2025). (Nanda et al., 2024) propose the use of inference-time optimization on sparse autoencoder decoder matrices to quickly evaluate and compare them.